OpenStudy (anonymous):

what is the base/original function of y=-2sqrt3x-12 -5?

OpenStudy (anonymous):

\[y=-2\sqrt{3x-12}-5\]

OpenStudy (paxpolaris):

\(y=\sqrt x\), ...or maybe... \(y=-\sqrt x\)

OpenStudy (anonymous):

thank you! I was also asked to describe the series of transformations from the base function to \[y=-2\sqrt{3x-12}-5\] do you have any ideas with the answer?

OpenStudy (paxpolaris):

\[y=-2\sqrt{3\left( x-4 \right)}-5\] if you start with \(y=\sqrt x\) > reflect across the x-axis, to get: \(y=\color{green}-\sqrt x\) > stretch vertically by 2: \(y=-\color{green}2\sqrt x\) > shrink horizontally by 3: \(y=-2\sqrt {\color{green}3x}\) > translate right by 4: \(y=-2\sqrt {3(x\color{green}{-4})}=-2\sqrt {3x-12}\) > translate down by 5: \(y==-2\sqrt {3x-12}\color{green}{-5}\)

OpenStudy (anonymous):

wow this is so helpful! Can you guide me through writing this equation in the form of \[y=af(k(x-c))+d\]?

OpenStudy (anonymous):

@PaxPolaris

OpenStudy (paxpolaris):

start with \(y=f(x)\) > if and only if \(a\) is negative, reflect across the x-axis, to get: \(y=-f(x)\) > stretch vertically by |a| : \(y=af(x)\) > if and only if \(k\) is negative, reflect across the y-axis: \(y=af(-x)\) > shrink horizontally by k : \(y=af(kx)\) > translate (move) horizontally by +c: \(y=af\left(k\left(x-c\right)\right)\) > translate (move) vertically by +d: \(y=af\left(k\left(x-c\right)\right)+d\)

OpenStudy (paxpolaris):

always do translations after you do stretch/shrink/reflect

OpenStudy (anonymous):

for the y=af(k(x-c))+d part, I'm still not quiet sure with the steps, can you show me the numbers and variables from the given function?

OpenStudy (anonymous):

@PaxPolaris

OpenStudy (paxpolaris):

you mean in:\[y=-2\sqrt{3\left( x-4 \right)}-5\] \[f(x)=\sqrt x\]\[a=-2\]\[k=3\]\[c=4\]\[d=-5\]

OpenStudy (anonymous):

yep, how do you write \[y=-2\sqrt{3x-1}\] in the function notation y=af(k(x-c))+d?

OpenStudy (anonymous):

@PaxPolaris

OpenStudy (paxpolaris):

\[=-2\sqrt{3\left( x-\frac13 \right)}\] \[f(x)=\sqrt x\]\[a=-2\]\[k=3\]\[c=1/3\]\[d=0\]

OpenStudy (anonymous):

what about y = -2f[3(x - 4)] - 5 ?

OpenStudy (anonymous):

@PaxPolaris

OpenStudy (paxpolaris):

that is what we did for the original question

OpenStudy (paxpolaris):

\[y=-2\sqrt{3\left( x-4 \right)}-5\]